OpenStudy (anonymous):

SOLVE FOR X: 24x ( 2+3) =

OpenStudy (anonymous):

120x

OpenStudy (anonymous):

120x

jhonyy9 (jhonyy9):

if 24x(2+3)=0 than 120x=0 --- x=0

OpenStudy (anonymous):

---x=0 i dnt undersstand this one

jhonyy9 (jhonyy9):

120x=0 x=0/120 x=0 - now is it clearly ?

OpenStudy (anonymous):

yes you are right

jhonyy9 (jhonyy9):

ok if it is right please you push the good answer button thank you